# Provable Benefit of Curriculum in Transformer Tree-Reasoning Post-Training

### v1

@@ Appendix D Details and Proofs of Autoregressive Reasoning Tree
D.1 Proof of Representation Theorem
@@
so that $\mathbf{V}\,\mathbf{E}[u]=\bm{\mu}^{u}$ (content only) for any token $u$, and attention logits depend only on positional encodings: $\mathbf{E}[u]^{\top}\mathbf{K}^{\top}\mathbf{Q}\,\mathbf{E}[z_{l-1}]=\mathbf{p}_{k}^{\top}\mathbf{W}\,\mathbf{p}_{d+1+l}$ if $u$ is at position $k$. Moreover, for each $l$ choose constants $c_{l}>0$ and impose a legality mask that sets logits of all $k\notin\mathcal{I}_{l}(z_{<l})$ to -$\infty$, while ensuring
@@
D.2 Proof of Finetuning Algorithms
@@
$\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}:=\sum_{j\in\mathcal{I}_{l}}\alpha_{l}(j)\,\bm{\mu}_{x_{j}},\qquad\ell_{r}:=\frac{\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle}{\beta},\ \ r\in\{0,...,\mathrm{EOS}\}.$
@@
$\pi_{\mathbf{W}}({i}_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}):=p_{\mathrm{vocab}}\big(u=\bm{\mu}_{x_{{i}_{l}}}\ \big|\ \hat{\bm{\mu}}^{\mathrm{att}}_{l+1}\big).$
@@
$\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}}({i}_{l}\mid\cdot)=\sum_{k\in\mathcal{I}_{l}}\alpha_{l}(k)\,\eta_{l}(k)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top},\quad\eta_{l}(k):=\frac{\big\langle\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \bm{\mu}_{x_{{i}_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\bm{\mu}_{r}\big\rangle}{\beta}.$
@@
$p_{\mathrm{vocab}}(r)=\frac{\exp(\ell_{r})}{\sum_{q}\exp(\ell_{q})},\qquad\ell_{r}=\frac{\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle}{\beta},$
@@
$\pi_{\mathbf{W}}({i}_{l}\mid\cdot)=p_{\mathrm{vocab}}(x_{{i}_{l}}):=p_{\mathrm{vocab}}\big(u=\bm{\mu}_{x_{{i}_{l}}}\mid\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}\big).$
We differentiate $\log p_{\mathrm{vocab}}(x_{{i}_{l}})$ w.r.t. $\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}$:
$\nabla_{\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}}\log p_{\mathrm{vocab}}(x_{{i}_{l}})$ $\stackrel{{\scriptstyle(1.1)}}{{=}}\nabla_{\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}}\Big(\ell_{x_{{i}_{l}}}-\log\sum_{r}e^{\ell_{r}}\Big)$ (13)
$\stackrel{{\scriptstyle(1.2)}}{{=}}\frac{\bm{\mu}_{x_{{i}_{l}}}}{\beta}-\sum_{r}\frac{e^{\ell_{r}}}{\sum_{q}e^{\ell_{q}}}\,\frac{\bm{\mu}_{r}}{\beta}$ (14)
$\stackrel{{\scriptstyle(1.3)}}{{=}}\frac{\bm{\mu}_{x_{{i}_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\,\bm{\mu}_{r}}{\beta}\ :=:\ \mathbf{g}_{l}^{\mathrm{vocab}}.$ (15)
Step (1.1) expands the log-softmax; (1.2) uses $\partial\ell_{r}/\partial\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}=\bm{\mu}_{r}/\beta$; (1.3) recognizes $p_{\mathrm{vocab}}$.
@@
$\nabla_{\mathbf{W}}\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}$ $=\nabla_{\mathbf{W}}\Big(\sum_{j\in\mathcal{I}_{l}}\alpha_{l}(j)\bm{\mu}_{x_{j}}\Big)$ (16)
$\stackrel{{\scriptstyle(2.1)}}{{=}}\sum_{j}\bm{\mu}_{x_{j}}\sum_{k}\frac{\partial\alpha_{l}(j)}{\partial s_{l}(k)}\,\nabla_{\mathbf{W}}s_{l}(k)$ (17)
$\stackrel{{\scriptstyle(2.2)}}{{=}}\sum_{j}\bm{\mu}_{x_{j}}\sum_{k}\alpha_{l}(j)\big(\delta_{jk}-\alpha_{l}(k)\big)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}$ (18)
$\stackrel{{\scriptstyle(2.3)}}{{=}}\sum_{k}\alpha_{l}(k)\Big(\bm{\mu}_{x_{k}}-\sum_{j}\alpha_{l}(j)\bm{\mu}_{x_{j}}\Big)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}$ (19)
$\stackrel{{\scriptstyle(2.4)}}{{=}}\sum_{k}\alpha_{l}(k)\big(\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}\big)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}.$ (20)
Step (2.1) is the chain rule; (2.2) uses the softmax Jacobian and $\nabla_{\mathbf{W}}s_{l}$; (2.3) collects terms; (2.4) recognizes $\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}$.
@@
$\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}}({i}_{l}\mid\cdot)$ $\stackrel{{\scriptstyle(3.1)}}{{=}}\Big(\nabla_{\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}}\log p_{\mathrm{vocab}}(x_{{i}_{l}})\Big)^{\top}\cdot\nabla_{\mathbf{W}}\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}$ (21)
$\stackrel{{\scriptstyle(3.2)}}{{=}}\sum_{k\in\mathcal{I}_{l}}\alpha_{l}(k)\,\underbrace{\Big\langle\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\ \mathbf{g}_{l}^{\mathrm{vocab}}\Big\rangle}_{\displaystyle:=\,\eta_{l}(k)}\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}.$ (22)
This yields the claimed positional outer-product decomposition with weights $\alpha_{l}(k)\eta_{l}(k)$, where $\mathbf{g}_{l}^{\mathrm{vocab}}=(\bm{\mu}_{x_{{i}_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\bm{\mu}_{r})/\beta$.
(4) Interchange and smoothness conditions. The above steps use: (i) $m_{l}$ is $\mathbf{W}$-independent, so $\nabla_{\mathbf{W}}s_{l}$ exists and is continuous; (ii) softmax is $C^{\infty}$, thus $\alpha_{l}$ and $p_{\mathrm{vocab}}$ are smooth in $\mathbf{W}$; (iii) boundedness of token embeddings $\bm{\mu}_{\cdot}$, vocabulary $\mathbf{U}$, and temperature $\beta>0$ ensures an $L^{1}$ dominator for Leibniz interchange when taking expectations over trajectories (used later in policy-gradient proofs).
@@
$\big|\eta_{l}(k)\big|\ \leq\ \frac{4}{\beta},\qquad\text{and}\qquad p_{\mathrm{vocab}}(r)=\mathrm{softmax}_{r}\big(\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle/\beta\big)$
@@
$\pi_{l}\ :=\ p_{\mathrm{vocab}}\big(u=\bm{\mu}_{x_{i_{l}}}\mid\hat{\bm{\mu}}^{\mathrm{att}}_{l+1}\big)\ \geq\ 1-\varepsilon_{l},\quad l=1,\ldots,k^{\star}+1.$
@@
Step 2 (vocabulary probability from attention concentration). With orthogonal embeddings, $\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{x_{i_{l}}}\rangle=\alpha_{l}(i_{l})$ and for any competitor token $r\neq x_{i_{l}}$, $\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle\leq 1-\alpha_{l}(i_{l})$ (competitors include tokens not present in the mixture, whose inner products are $0\leq 1-\alpha_{l}$). Hence the vocabulary-logit gap between the correct token and any competitor is at least
@@
$\pi_{l}\ =\ \frac{e^{\alpha_{l}(i_{l})/\beta}}{\sum_{r}e^{\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle/\beta}}\ \geq\ \frac{1}{1+\sum_{r\neq x_{i_{l}}}e^{-\gamma_{l}^{\mathrm{vocab}}/\beta}}\ \geq\ \frac{1}{1+M\,e^{-(2\alpha_{l}(i_{l})-1)/\beta}}.$
@@
Let the input $\mathbf{x}$ be fixed and the decoding length be $k^{\star}\geq 2$. Denote the generated token sequence by $\hat{\bm{\mu}}^{z_{1:k^{\star}}}=(\hat{\bm{\mu}}^{z_{1}},\dots,\hat{\bm{\mu}}^{z_{k^{\star}}})$. At step $l$ ($1\leq l\leq k^{\star}+1$), an attention policy first samples a secret index’s token ${i}_{l}\sim\pi_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}):=\hat{p}_{\mathbf{\mathbf{W}}}\big(u=\bm{\mu}_{x_{{i}_{l}}}\,\big|\,\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}\big)$. The next token is then deterministically produced by the deterministic Feedforward $\operatorname{FFN}_{m}(\cdot)$ at $m$-th reasoning procedure, so that the only source of randomness is the sampling of tokens $\bm{\mu}_{x_{{i}_{l}}}$ corresponding to secret index sequences ${i}_{1:k^{\star}+1}$. We therefore define the (random) trajectory as $\tau:=({i}_{1:k^{\star}+1})$ and the induced measure
$p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})=\prod_{l=1}^{k^{\star}+1}\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}\big),$
where the tokens $\hat{\bm{\mu}}^{z_{1:k^{\star}}}$ are deterministic functions of $\tau$ via $g_{\mathbf{W}^{k^{\star}}}$. Let the terminal reward be $R^{k^{\star}}(\operatorname{TF}(\cdot;\mathbf{W}))$, which does not explicitly depend on $\mathbf{W}^{k^{\star}}$. Define the population loss
$\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}(\mathbf{W}^{k^{\star}})=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}_{x},\tau\sim p_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x})}\Big[R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\Big].$ (23)
@@
$\nabla_{\mathbf{W}^{k^{\star}}}\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}(\mathbf{W}^{k^{\star}})=\mathbb{E}_{\tau}\Bigg[R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\cdot\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\bm{\mu}}^{z_{1:l}}\big)\Bigg].$ (24)
@@
$\begin{aligned} &\nabla_{\mathbf{W}^{k^{\star}}}\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}(\mathbf{W}^{k^{\star}})=\mathbb{E}_{\tau}\Bigg[R^{k^{\star}}\big(\hat{\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\sum_{l=1}^{k^{\star}+1}\sum_{k\in\mathcal{I}_{l}}\alpha^{k^{\star}}_{l}(k)\,\eta^{k^{\star}}_{l}(k)\;\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}\Bigg],\end{aligned}$
@@
  • $\hat{\bm{\mu}}_{l+1}^{\mathrm{att},k^{\star}}:=\sum_{j\in\mathcal{I}_{l}}\alpha_{l}^{k^{\star}}(j)\bm{\mu}_{x_{j}}$,
  • $\eta_{l}^{k^{\star}}(k):=\langle\bm{\mu}_{x_{k}}-\hat{\bm{\mu}}_{l+1}^{\mathrm{att},k^{\star}},\bm{\mu}_{x_{{i}_{l}}}-\sum_{r}p^{k^{\star}}_{\mathrm{vocab}}(r)\bm{\mu}_{r}\rangle/\beta$,
  • $p^{k^{\star}}_{\mathrm{vocab}}(\cdot):=\frac{\exp(\frac{\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{r}\rangle}{\beta})}{\sum_{q}\exp(\frac{\langle\hat{\bm{\mu}}^{\mathrm{att}}_{l+1},\bm{\mu}_{q}\rangle}{\beta})}$.
Proof.Write $\tau=({i}_{1:k^{\star}+1})$ and abbreviate $\mathbb{E}_{\tau}$ for $\mathbb{E}_{\tau\sim p_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x})}$. The tokens $\hat{\bm{\mu}}^{z_{1:k^{\star}}}(\tau)$ are deterministic given $\tau$ via $g_{\mathbf{W}^{k^{\star}}}$.

### v2

@@ Appendix G Details and Proofs of Autoregressive Reasoning Tree
G.1 Proof of Representation Theorem
@@
so that $\mathbf{V}\,\mathbf{E}[x_{m}]=p_{m}$ (position only) for any token $x_{m}$, and attention logits depend only on positional encodings: $\mathbf{E}[u]^{\top}\mathbf{K}^{\top}\mathbf{Q}\,\mathbf{E}[z_{l-1}]=\mathbf{p}_{k}^{\top}\mathbf{W}\,\mathbf{p}_{d+1+l}$, $\mathbf{p}_{d+1+l}:=\mathbf{p}_{i_{l}}$ if $u$ is at position $k$. Moreover, for each $l$ choose constants $c_{l}>0$ and impose a legality mask that sets logits of all $k\notin\mathcal{I}_{l}(z_{<l})$ to -$\infty$, while ensuring
@@
G.2 Proof of Finetuning Algorithms
@@
$\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}:=\sum_{j\in\mathcal{I}_{l}}\alpha_{l}(j)\,\mathbf{p}_{j},\qquad\ell_{r}:=\frac{\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle}{\beta},\ \ r\in\{1,...,d+1\}.$
@@
$\pi_{\mathbf{W}}({i}_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}):=p_{\mathrm{vocab}}\big(u=\bm{\mu}_{x_{{i}_{l}}}\ \big|\ \hat{\mathbf{p}}^{\mathrm{att}}_{l+1}\big).$
@@
$\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}}({i}_{l}\mid\cdot)=\sum_{k\in\mathcal{I}_{l}}\alpha_{l}(k)\,\eta_{l}(k)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top},\quad\eta_{l}(k):=\frac{\big\langle\mathbf{p}_{k}-\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \mathbf{p}_{{i}_{l}}-\sum_{r}p_{\mathrm{vocab}}(r)\mathbf{p}_{r}\big\rangle}{\beta}.$
@@
$p_{\mathrm{vocab}}(r)=\frac{\exp(\ell_{r})}{\sum_{q}\exp(\ell_{q})},\qquad\ell_{r}=\frac{\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle}{\beta},$
@@
$\pi_{\mathbf{W}}({i}_{l}\mid\cdot)=p_{\mathrm{vocab}}(x_{{i}_{l}}):=p_{\mathrm{vocab}}\big(u=\bm{\mu}_{x_{{i}_{l}}}\mid\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}\big).$
We differentiate $\log p_{\mathrm{vocab}}(x_{{i}_{l}})$ w.r.t. $\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}$:
$\nabla_{\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}}\log p_{\mathrm{vocab}}(x_{{i}_{l}})$ $\stackrel{{\scriptstyle(1.1)}}{{=}}\nabla_{\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}}\Big(\ell_{x_{{i}_{l}}}-\log\sum_{r}e^{\ell_{r}}\Big)$ (13)
$\stackrel{{\scriptstyle(1.2)}}{{=}}\frac{\mathbf{p}_{{i}_{l}}}{\beta}-\sum_{r}\frac{e^{\ell_{r}}}{\sum_{q}e^{\ell_{q}}}\,\frac{\mathbf{p}_{r}}{\beta}$ (14)
$\stackrel{{\scriptstyle(1.3)}}{{=}}\frac{\mathbf{p}_{{i}_{l}}-\sum_{r}p_{\mathrm{vocab}}(r)\,\mathbf{p}_{r}}{\beta}\ :=:\ \mathbf{g}_{l}^{\mathrm{vocab}}.$ (15)
Step (1.1) expands the log-softmax; (1.2) uses $\partial\ell_{r}/\partial\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}=\mathbf{p}_{r}/\beta$; (1.3) recognizes $p_{\mathrm{vocab}}$.
@@
$\nabla_{\mathbf{W}}\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}$ $=\nabla_{\mathbf{W}}\Big(\sum_{j\in\mathcal{I}_{l}}\alpha_{l}(j)\mathbf{p}_{j}\Big)$ (16)
$\stackrel{{\scriptstyle(2.1)}}{{=}}\sum_{j}\mathbf{p}_{j}\sum_{k}\frac{\partial\alpha_{l}(j)}{\partial s_{l}(k)}\,\nabla_{\mathbf{W}}s_{l}(k)$ (17)
$\stackrel{{\scriptstyle(2.2)}}{{=}}\sum_{j}\mathbf{p}_{j}\sum_{k}\alpha_{l}(j)\big(\delta_{jk}-\alpha_{l}(k)\big)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}$ (18)
$\stackrel{{\scriptstyle(2.3)}}{{=}}\sum_{k}\alpha_{l}(k)\Big(\mathbf{p}_{x_{k}}-\sum_{j}\alpha_{l}(j)\mathbf{p}_{j}\Big)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}$ (19)
$\stackrel{{\scriptstyle(2.4)}}{{=}}\sum_{k}\alpha_{l}(k)\big(\mathbf{p}_{k}-\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}\big)\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}.$ (20)
Step (2.1) is the chain rule; (2.2) uses the softmax Jacobian and $\nabla_{\mathbf{W}}s_{l}$; (2.3) collects terms; (2.4) recognizes $\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}$.
@@
$\nabla_{\mathbf{W}}\log\pi_{\mathbf{W}}({i}_{l}\mid\cdot)$ $\stackrel{{\scriptstyle(3.1)}}{{=}}\Big(\nabla_{\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}}\log p_{\mathrm{vocab}}(x_{{i}_{l}})\Big)^{\top}\cdot\nabla_{\mathbf{W}}\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}$ (21)
$\stackrel{{\scriptstyle(3.2)}}{{=}}\sum_{k\in\mathcal{I}_{l}}\alpha_{l}(k)\,\underbrace{\Big\langle\mathbf{p}_{k}-\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\ \mathbf{g}_{l}^{\mathrm{vocab}}\Big\rangle}_{\displaystyle:=\,\eta_{l}(k)}\,\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}.$ (22)
This yields the claimed positional outer-product decomposition with weights $\alpha_{l}(k)\eta_{l}(k)$, where $\mathbf{g}_{l}^{\mathrm{vocab}}=(\mathbf{p}_{{{i}_{l}}}-\sum_{r}p_{\mathrm{vocab}}(r)\mathbf{p}_{r})/\beta$.
(4) Interchange and smoothness conditions. The above steps use: (i) $m_{l}$ is $\mathbf{W}$-independent, so $\nabla_{\mathbf{W}}s_{l}$ exists and is continuous; (ii) softmax is $C^{\infty}$, thus $\alpha_{l}$ and $p_{\mathrm{vocab}}$ are smooth in $\mathbf{W}$; (iii) boundedness of token embeddings $\mathbf{p}_{\cdot}$, vocabulary $\mathbf{U}$, and temperature $\beta>0$ ensures an $L^{1}$ dominator for Leibniz interchange when taking expectations over trajectories (used later in policy-gradient proofs).
@@
$\big|\eta_{l}(k)\big|\ \leq\ \frac{4}{\beta},\qquad\text{and}\qquad p_{\mathrm{vocab}}(r)=\mathrm{softmax}_{r}\big(\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle/\beta\big)$
@@
$\pi_{l}\ :=\ p_{\mathrm{vocab}}\big(u=\mathbf{p}_{{i_{l}}}\mid\hat{\mathbf{p}}^{\mathrm{att}}_{l+1}\big)\ \geq\ 1-\varepsilon_{l},\quad l=1,\ldots,k^{\star}+1.$
@@
Step 2 (vocabulary probability from attention concentration). With orthogonal embeddings, $\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{{i_{l}}}\rangle=\alpha_{l}(i_{l})$ and for any competitor token $r\neq x_{i_{l}}$, $\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle\leq 1-\alpha_{l}(i_{l})$ (competitors include tokens not present in the mixture, whose inner products are $0\leq 1-\alpha_{l}$). Hence the vocabulary-logit gap between the correct token and any competitor is at least
@@
$\pi_{l}\ =\ \frac{e^{\alpha_{l}(i_{l})/\beta}}{\sum_{r}e^{\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle/\beta}}\ \geq\ \frac{1}{1+\sum_{r\neq x_{i_{l}}}e^{-\gamma_{l}^{\mathrm{vocab}}/\beta}}\ \geq\ \frac{1}{1+M\,e^{-(2\alpha_{l}(i_{l})-1)/\beta}}.$
@@
Let the input $\mathbf{x}$ be fixed and the decoding length be $k^{\star}\geq 2$. Denote the generated token sequence by $\hat{\mathbf{p}}^{z_{1:k^{\star}}}=(\hat{\mathbf{p}}^{z_{1}},\dots,\hat{\mathbf{p}}^{z_{k^{\star}}})$. At step $l$ ($1\leq l\leq k^{\star}+1$), an attention policy first samples a secret index’s token ${i}_{l}\sim\pi_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}):=\hat{p}_{\mathbf{\mathbf{W}}}\big(u=\bm{\mu}_{x_{{i}_{l}}}\,\big|\,\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}\big)$. The next token is then deterministically produced by the deterministic Feedforward $\operatorname{FFN}_{m}(\cdot)$ at $m$-th reasoning procedure, so that the only source of randomness is the sampling of tokens $\bm{\mu}_{x_{{i}_{l}}}$ corresponding to secret index sequences ${i}_{1:k^{\star}+1}$. We therefore define the (random) trajectory as $\tau:=({i}_{1:k^{\star}+1})$ and the induced measure
$p_{\mathbf{W}^{k^{\star}}}(\tau\mid\mathbf{x})=\prod_{l=1}^{k^{\star}+1}\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}\big),$
where the tokens $\hat{\mathbf{p}}^{z_{1:k^{\star}}}$ are deterministic functions of $\tau$ via $g_{\mathbf{W}^{k^{\star}}}$. Let the terminal reward be $R^{k^{\star}}(\operatorname{TF}(\cdot;\mathbf{W}))$, which does not explicitly depend on $\mathbf{W}^{k^{\star}}$. Define the population loss
$\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}(\mathbf{W}^{k^{\star}})=\mathbb{E}_{\mathbf{x}\sim\mathcal{P}_{x},\tau\sim p_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x})}\Big[R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\Big].$ (23)
@@
$\nabla_{\mathbf{W}^{k^{\star}}}\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}(\mathbf{W}^{k^{\star}})=\mathbb{E}_{\tau}\Bigg[R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\cdot\sum_{l=1}^{k^{\star}+1}\nabla_{\mathbf{W}^{k^{\star}}}\log\pi_{\mathbf{W}^{k^{\star}}}\big({i}_{l}\mid\mathbf{x},\hat{\mathbf{p}}^{z_{1:l}}\big)\Bigg].$ (24)
@@
$\begin{aligned} &\nabla_{\mathbf{W}^{k^{\star}}}\mathcal{J}_{\mathrm{REINFORCE}}^{k^{\star}}(\mathbf{W}^{k^{\star}})=\mathbb{E}_{\tau}\Bigg[R^{k^{\star}}\big({\bm{\mu}}^{z_{k^{\star}}}(\tau_{-2})\big)\sum_{l=1}^{k^{\star}+1}\sum_{k\in\mathcal{I}_{l}}\alpha^{k^{\star}}_{l}(k)\,\eta^{k^{\star}}_{l}(k)\;\mathbf{p}_{k}\mathbf{p}_{c_{l}}^{\top}\Bigg],\end{aligned}$
@@
  • $\hat{\mathbf{p}}_{l+1}^{\mathrm{att},k^{\star}}:=\sum_{j\in\mathcal{I}_{l}}\alpha_{l}^{k^{\star}}(j)\mathbf{p}_{j}$,
  • $\eta_{l}^{k^{\star}}(k):=\langle\mathbf{p}_{k}-\hat{\mathbf{p}}_{l+1}^{\mathrm{att},k^{\star}},\mathbf{p}_{{{i}_{l}}}-\sum_{r}p^{k^{\star}}_{\mathrm{vocab}}(r)\mathbf{p}_{r}\rangle/\beta$,
  • $p^{k^{\star}}_{\mathrm{vocab}}(\cdot):=\frac{\exp(\frac{\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{r}\rangle}{\beta})}{\sum_{q}\exp(\frac{\langle\hat{\mathbf{p}}^{\mathrm{att}}_{l+1},\mathbf{p}_{q}\rangle}{\beta})}$.
Proof.Write $\tau=({i}_{1:k^{\star}+1})$ and abbreviate $\mathbb{E}_{\tau}$ for $\mathbb{E}_{\tau\sim p_{\mathbf{W}^{k^{\star}}}(\cdot\mid\mathbf{x})}$. The tokens $\hat{\mathbf{p}}^{z_{1:k^{\star}}}(\tau)$ are deterministic given $\tau$ via $g_{\mathbf{W}^{k^{\star}}}$.